# Unified Optimization of Source Weights and Transfer Quantities in Multi-Source Transfer Learning: An Asymptotic Framework

### v1

@@ V-A Experimental Settings
2:Parameter: Learning rate $\eta$.
@@
4:Output: a well-trained $\underline{\theta}_{0}$ for target task model $f_{{\underline{\theta}}_{0}}$ .
@@
1:Input: Datasets $\{D_{S_{k}}=\{(z_{S_{k}}^{j},y_{S_{k}}^{j})\}_{j=1}^{N_{k}}\}_{k=1}^{K}$, model type $f_{\underline{\theta}}$, parameters $\{\underline{\theta}_{k}\}_{k=1}^{K}$, parameter dimension $d$, weights $\{w_{k}\}_{k=1}^{K}$
@@
3:Initialize: Randomly initialize $\{\underline{\theta}_{k}\}_{k=1}^{K}$, set weights $w_{k^{\prime}}^{(k)}=0$ for all $k\neq k^{\prime}$
@@
7:    $\mathcal{L}_{train}^{(k)}\leftarrow\frac{,\sum\limits_{(z^{j},y^{j})\in{D_{S_{k}}}}\ell\left(y^{j},f_{\underline{\theta}_{k}}(z^{j})\right)}{|\{D_{S_{k}}\}_{k=1}^{K}|}$    $\frac{\sum\limits_{k^{\prime}\neq k}\sum\limits_{(z^{j},y^{j})\in{D_{S_{k^{\prime}}}}}\ell\left(y^{j},f_{\underline{\theta}_{k}}(z^{j})\right)\cdot w_{k^{\prime}}^{(k)}}{|\{D_{S_{k}}\}_{k=1}^{K}|}$
@@
Figure 3: Overview of the Datasets: Examples from the cross-domain datasets DomainNet and OfficeHome, where images from different domains exhibit distinct visual styles or are captured using different devices.
In the multi-source transfer learning setting, our goal is to transfer useful knowledge from multiple source domains to a specific target domain. In particular, our evaluation is performed in the 10-shot learning setting, where only 10 labeled samples per class are available in the target domain.
In multi-task learning setting, every task is simultaneously a target task to be learned and a source task from which knowledge is transferred to other tasks. Specifically, our approach iteratively selects one task as the target while treating others as sources, and derives weights through our theory. For each task, we construct its training set by combining all samples from the target domain with the weighted samples from the remaining domains, where the weights are again computed by UOWQ. All domain-specific tasks are then jointly optimized in a unified multi-task learning framework.
@@ C I P Q R S Avg Ar Cl Pr Rw Avg
TABLE IV: Multi-Task Learning Performance on DomainNet and Office-Home. Each column indicates the target domain used for evaluation. The highest and second-highest accuracy are marked in bold and underline, respectively.
Method DomainNet Office-Home
C I P Q R S Avg Ar Cl Pr Rw Avg
EW 67.3 22.7 56.1 35.1 69.3 55.9 51.1 63.0 76.5 88.5 77.7 76.4
MGDA-UB[41] 69.0 22.4 56.1 31.7 69.7 55.9 50.8 64.3 75.3 89.7 79.3 77.2
GradNorm[7] 67.1 22.5 56.1 35.7 69.3 55.8 51.1 65.5 75.3 88.7 78.9 77.1
PCGrad[57] 67.8 22.7 56.0 34.7 69.2 56.1 51.1 63.9 76.0 88.9 78.3 76.8
CAGrad[27] 68.4 22.7 55.7 32.1 69.8 55.6 50.7 63.8 75.9 89.1 78.3 76.8
RGW[26] 67.1 21.9 55.2 33.9 68.6 55.6 50.4 65.1 78.7 88.7 79.9 78.1
MoCo[13] 60.3 19.0 46.6 39.9 57.9 50.7 45.7 64.1 79.8 89.6 79.6 78.3
MoDo[3] 67.9 22.0 55.4 36.0 69.4 56.3 51.2 66.2 78.2 89.8 80.3 78.7
UOWQ 68.2 22.7 56.5 37.2 70.0 57.2 51.9 69.3 77.3 89.5 80.4 79.1
V-B Experimental Design and Model Adaptation
For the multi-source transfer learning setting, we follow a unified evaluation setup for all methods. The baselines include: 1) Model-Weighting Based Few-Shot Methods: H-ensemble [55], MCW [21]. 2) Sample-Based Few-Shot Methods: MADA [61], WADN [44] 3) Source Ablation Studies: Target-Only, Single-Source-Avg/Single-Source-Best (average/best single-source transfer results).
V-C Few-Shot Multi-Source Transfer Performance
Sample Weighting vs. Model Weighting. In the multi-source transfer learning setting, methods based on sample weighting consistently exhibit stronger performance than those relying on model-level aggregation. Specifically, approaches that directly optimize a mixture of source and target samples—such as WADN, MADA, OTQMS, and UOWQ—tend to outperform model-weighting techniques like H-ensemble and MCW, which combine predictions or parameters from multiple pretrained models. A plausible explanation is that sample-weighting strategies allow the target model to interact directly with source data during training, enabling more flexible adaptation to target-specific distributions, whereas model-weighting methods are constrained by fixed source representations. Among sample-weighting approaches, UOWQ further improves upon existing baselines by jointly optimizing both the amount of transferred data and the corresponding weights, whereas OTQMS focuses on optimizing transfer quantities and WADN primarily adjusts weighting coefficients.
Gains over Strong Baselines in Challenging Domains. Beyond average accuracy improvements, UOWQ consistently outperforms strong baselines in the most challenging domains. For example, in the experiments of DomainNet dataset , UOWQ improves over the previous best results in the Quickdraw target domain by 3.0% absolute accuracy (36.8% vs. 33.8%) and in the Painting target domain by 1.4% (35.2% vs. 33.8%). These gains in visually and semantically divergent domains demonstrate that UOWQ’s sample-weighting strategy is particularly effective in mitigating domain shift.
V-D Multi-Task Learning Performance
The results for the multi-task learning setting are summarized in Table IV. We make the following observations.
Theoretical Consistency. Our theoretical analysis provides a clear explanation for the empirical observations in Table IV. Specifically, Theorems 5 shows that using the theoretically derived task weights leads to more balanced contributions from different tasks during multi-task optimization. This balanced weighting reduces the risk of over-emphasizing a single task and helps alleviate negative transfer caused by task mismatch. Empirically, this theoretical prediction is reflected in the experimental results: UOWQ achieves stable performance improvements across multiple target domains without exhibiting noticeable performance trade-offs. Consequently, the results on DomainNet and Office-Home provide empirical support for the effectiveness of the proposed weighting theory in multi-task learning settings.
V-E Weight Visualization
To analyze the domain preference learned by UOWQ, we visualize the coefficients $\alpha$ obtained from the quadratic optimization problem together with the base weights $w$ in Fig. 4, where each row corresponds to a target domain and each column corresponds to a source domain. The visualizations reveal clear and consistent patterns across both benchmarks. On DomainNet, when Art, Clipart, and Product are treated as target domains, the real-world domain consistently receives the largest $\alpha$ coefficients. A similar trend is observed on Office-Home for Art and Clipart targets. Overall, these results indicate that the quadratic optimization naturally allocates higher transfer coefficients to more transferable source domains, consistent with prior empirical observations in [25].
(a) DomainNet: $\alpha$ weights
(b) DomainNet: $w$ weights
(c) Office-Home: $\alpha$ weights
(d) Office-Home: $w$ weights
V-F Robustness under Different Shot Settings
We further investigate the robustness of UOWQ with respect to the availability of target samples by gradually increasing the number of shots per class from 5 to 100. Fig. 5 reports the results on the Office-Home dataset across all four target domains (Ar, Cl, Pr, and Rw), in comparison with Target-Only and AllSources $\cup$ Target. Across all domains, UOWQ maintains consistently competitive accuracy throughout the entire range of shot settings, showing neither performance collapse in low-shot regimes nor saturation-induced degradation as more target samples are introduced. In contrast to methods that rely heavily on either source aggregation or target-only supervision, UOWQ exhibits stable and monotonic performance trends as the number of shots increases.ă These observations indicate that the proposed approach adapts smoothly to different data regimes and remains robust under varying levels of target supervision.
(a) Ar
(b) Cl
(c) Pr
(d) Rw
V-G Static vs. Dynamic Weighting.
@@ V-F Robustness under Different Shot Settings
V-H Computational Efficiency.
We further compare the computational cost of different methods on the DomainNet benchmark, and report the wall-clock training time in Table VI. Compared with MADA, which requires substantially longer training time across all target domains, both AllSources $\cup$ Target and UOWQ exhibit significantly improved efficiency. Notably, UOWQ achieves comparable training time to AllSources $\cup$ Target, with only a marginal overhead introduced by the additional optimization procedure. Across all domains, the average training time of UOWQ (10:00:32) remains close to that of AllSources $\cup$ Target (09:50:10), and is substantially lower than MADA (17:07:55). These results indicate that the proposed method introduces minimal computational overhead while providing consistent performance gains, demonstrating a favorable trade-off between effectiveness and efficiency.
TABLE VI: Training time comparison on the DomainNet dataset across different target domains.
Method Backbone DomainNet
[…]
V-I Compatibility to LoRA-Based Training
We further examine the compatibility of UOWQ with LoRA. Specifically, we apply UOWQ to multi-source transfer on the Office-Home dataset using the ViT-B backbone. Each domain is treated as a distinct source task, and LoRA modules with rank 8 are inserted into the projection layers of transformer blocks. As summarized in Table VII, UOWQ consistently improves accuracy over baselines, confirming that the proposed theoretical principles remain effective under LoRA-based adaptation.
TABLE VII: Multi-Source Transfer with LoRA on Office-Home. We apply LoRA on ViT-B backbone for PEFT.
@@ V-I Compatibility to LoRA-Based Training
VI Conclusion
@@
In this work, we propose UOWQ, a theoretical framework for the unified optimization of source weights and transfer quantities in multi-source transfer learning. By formulating the training process as a parameter estimation problem and conducting an asymptotic analysis of a Kullback–Leibler–based generalization error measure, we derived a principled method for computing the optimal source weights and transfer quantities. A notable theoretical insight is that, once the weights are jointly optimized, the optimal solution always utilizes all source samples, while the adaptive weighting mechanism effectively suppresses the bias introduced by heterogeneous domains.
Building on these theoretical results, we further developed a practical algorithm that assigns sample-level weights and updates them dynamically during training. This design enables the framework to adapt to evolving model parameters and supports both multi-source transfer learning and multi-task learning scenarios.
Extensive experiments on real-world datasets demonstrate that UOWQ consistently improves performance over state-of-the-art baselines, particularly in few-shot and cross-domain settings. These results validate both the theoretical foundations and the practical advantages of our approach.
Appendix A Proof of Theorem 2
@@
In this setting, the MLE and its expected definition are given as follows.
@@
Combining with (38), we can get
@@
Applying the Central Limit Theorem to (A), we can get
$\sqrt{N_{0}+n_{1}}L_{n}^{{}^{\prime}}(E_{\hat{\theta}})\xrightarrow{a.s.}$ (41)
@@
By combining (41), (A), and (A), we can get
@@
Lemma 8.
[…]
$=\frac{1}{2}{J}(\theta_{0})\text{MSE}(\hat{\theta})+o(\frac{1}{N_{0}}),$ (50)
Proof.
$D\left(P_{X;\theta_{0}}\middle\|P_{X;\hat{\theta}(X^{N_{0}})}\right)$
$=\sum_{x\in X}P_{X;\theta_{0}}(x)\log\frac{P_{X;\theta_{0}}(x)}{P_{X;\hat{\theta}(X^{N_{0}})}(x)}$
We denote $\delta$ as a small constant, and we can rewrite (1) as
$=\sum_{X^{N_{0}}}P_{X^{n};\theta_{0}}(X^{N_{0}})\Bigg(\frac{1}{2}\sum_{x\in X}\frac{\left(P_{X;\hat{\theta}(X^{N_{0}})}(x)-P_{X;\theta_{0}}(x)\right)^{2}}{P_{X;\theta_{0}}(x)}$
$+o(|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2})\Bigg)$
$=\sum_{X^{N_{0}}}P_{X^{n};\theta_{0}}(X^{N_{0}})\Bigg(\frac{1}{2}\sum_{x\in X}\frac{\Bigg(\frac{\partial P_{X;\theta_{0}}(x)}{\partial\theta_{0}}(\hat{\theta}(X^{N_{0}})-\theta_{0})\Bigg)^{2}}{P_{X;\theta_{0}}(x)}+$
$o(|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2})\Bigg)$
$+\sum_{\left\{X^{N_{0}}:|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2}\geq\delta\right\}}P_{X^{n};\theta_{0}}(X^{N_{0}})\cdot$
Definition 9.
[…]
We denote $\hat{P}_{X^{N_{0}}}$ as the empirical distribution of $X^{N_{0}}$. Applying Sanov’s Theorem to (A), we can know that
Then, we aim to establish a connection between (54) and $|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2}$. From (A), we can know that the $D\left(\hat{P}_{X^{N_{0}}}\middle\|P_{X;\theta_{0}}\right)$ in (54) can be transformed to
$D\left(\hat{P}_{X^{N_{0}}}\middle\|P_{X;\theta_{0}}\right)$
From the characteristics of MLE, we can know that
$=0$
$+O(|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2}),$ (56)
which can be transformed to
Using the Cauchy-Schwarz inequality, we can obtain
where
Combining with (A), (A), and (A), the inequality (A) can be transformed to
$D\left(\hat{P}_{X^{N_{0}}}\middle\|P_{X;\theta_{0}}\right)$
$+o(|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2})$
Combining (54) and (A), we can know that
For the first term in (A), the $|\hat{\theta}(X^{N_{0}})-\theta_{0}|$ is small enough for us to omit the term $o(|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2})$. As for the second term in in (A), even though the magnitude of $|\hat{\theta}(X^{N_{0}})-\theta_{0}|$ is no longer negligible, the probability of such sequences is $O(e^{\frac{-N_{0}J(\theta_{0})|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2}}{2}})$ by (61), which is exponentially decaying with $N_{0}$ such that the second term is $o{(\frac{1}{N_{0}})}$. By transferring (A), we can get
$\left(\frac{1}{2}\sum_{x\in X}\frac{\left(\frac{\partial P_{X;\theta_{0}}(x)}{\partial\theta_{0}}(\hat{\theta}(X^{N_{0}})-\theta_{0})\right)^{2}}{P_{X;\theta_{0}}(x)}\right)+o(|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2})\cdot$
$+\sum_{\left\{X^{N_{0}}:|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2}\geq\delta\right\}}P_{X^{n};\theta_{0}}(X^{N_{0}})\cdot$
$\left(\frac{1}{2}\sum_{x\in X}\frac{\left(\frac{\partial P_{X;\theta_{0}}(x)}{\partial\theta_{0}}(\hat{\theta}(X^{N_{0}})-\theta_{0})\right)^{2}}{P_{X;\theta_{0}}(x)}\right)+o{(\frac{1}{N_{0}})}$
$=\frac{1}{2}\mathbb{E}\left[\sum_{x\in X}\frac{\left(\frac{\partial P_{X;\theta_{0}}(x)}{\partial\theta_{0}}(\hat{\theta}(X^{N_{0}})-\theta_{0})\right)^{2}}{P_{X;\theta_{0}}(x)}\right]+o{(\frac{1}{N_{0}})}.$ (62)
We then transform (A) with (A)
Combining (A), (A), we can get
$\mathbb{E}\left[D\left(P_{X;\theta_{0}}\middle\|P_{X;\hat{\theta}}\right)\right]=\frac{1}{2}\mathbb{E}\left[\left(\hat{\theta}-\theta_{0}\right)^{2}\right]{J}(\theta_{0})+o{(\frac{1}{N_{0}})}.$ (64)
Finally, we make a decomposition for the K-L measure
[…]
$=\frac{1}{2}{J}(\theta_{0})\left(\mathbb{E}\left[\left(\hat{\theta}-E_{\hat{\theta}}\right)^{2}\right]+\mathbb{E}\left[\left(E_{\hat{\theta}}-\theta_{0}\right)^{2}\right]\right)+o(\frac{1}{N_{0}}).$ (65)
Substituting (27) and (49) into (A), we have
@@
Similarly, by taking the derivative of (A) with respect to $n_{1}$, we observe that the derivative is
@@
Appendix B Proof of Proposition 3

### v2

@@ VI-A Experimental Settings
Figure 3: Overview of the Datasets: Examples from the cross-domain datasets DomainNet and Office-Home, where images from different domains exhibit distinct visual styles or are captured using different devices.
@@
2:Parameter: Learning rate $\eta$
@@
4:Output: a well-trained $\underline{\theta}_{0}$ for target task model $f_{{\underline{\theta}}_{0}}$
@@
1:Input: Datasets $\{D_{S_{k}}=\{(z_{S_{k}}^{j},y_{S_{k}}^{j})\}_{j=1}^{N_{k}}\}_{k=1}^{K}$, model type $f_{\underline{\theta}}$, parameters $\{\underline{\theta}_{k}\}_{k=1}^{K}$, parameter dimension $d$, weights $\{w_{k^{\prime}}^{(k)}\}_{k,k^{\prime}=1,\;k^{\prime}\neq k}^{K}$
@@
3:Initialize: Randomly initialize $\{\underline{\theta}_{k}\}_{k=1}^{K}$, set weights $w_{k^{\prime}}^{(k)}=0$ for all $k^{\prime}\neq k$
@@
7:    $\mathcal{L}_{train}^{(k)}\leftarrow\frac{\sum\limits_{(z^{j},y^{j})\in{D_{S_{k}}}}\ell\left(y^{j},f_{\underline{\theta}_{k}}(z^{j})\right)}{|\{D_{S_{k}}\}_{k=1}^{K}|}$    $\frac{\sum\limits_{k^{\prime}\neq k}\sum\limits_{(z^{j},y^{j})\in{D_{S_{k^{\prime}}}}}\ell\left(y^{j},f_{\underline{\theta}_{k}}(z^{j})\right)\cdot w_{k^{\prime}}^{(k)}}{|\{D_{S_{k}}\}_{k=1}^{K}|}$
@@
In the multi-source transfer learning setting, our goal is to transfer useful knowledge from multiple source domains to a specific target domain. ă In the experiments under this setting, our evaluation is performed in the supervised-10-shot learning setting, where only 10 labeled samples per class are available in the target domain, while all available samples from the source domains in the dataset are utilized.
In the multi-task learning setting, each task simultaneously serves as both a target task to be learned and a source task from which knowledge is transferred to other tasks. In the experiments under this setting, all available data from every domain in the dataset are fully utilized.
@@
Model-Weighting Based Few-Shot Methods:
MCW[22] ViT-S 54.9 21.0 53.6 20.4 70.8 42.4 43.9 68.9 48.0 77.4 86.0 70.1
Sample-Based Few-Shot Methods:
@@ C I P Q R S Avg Ar Cl Pr Rw Avg
TABLE IV: Training time comparison on the DomainNet dataset across different target domains.
Method Backbone DomainNet
[…]
@@
TABLE VI: Multi-Source Transfer with LoRA on Office-Home. We apply LoRA on ViT-B backbone for PEFT.
@@
ă
VI-B Experimental Design and Model Adaptation
For the multi-source transfer learning setting, we follow a unified evaluation setup for all methods. The baselines include: 1) Source Ablation Studies: Target-Only, Single-Source-Avg/Single-Source-Best (average/best single-source transfer results), AllSources $\cup$ Target (transfer results using all available source data without weighting).ă 2) Model-Weighting Based Few-Shot Methods: H-ensemble [55], MCW [22]. 3) Sample-Based Few-Shot Methods: MADA [62], WADN [44], OTQMS [60]. Furthermore, we report the results of several unsupervised methods for reference, including MSFDA [43], DATE [15], M3SDA [39]. For these unsupervised methods, we adopt a different experimental setting: instead of following the supervised-10-shot setting, we utilize all available unlabeled target data, corresponding to an unsupervised-all-shot setting.
VI-C Few-Shot Multi-Source Transfer Performance
Sample-Based vs. Model-Based. In the multi-source transfer learning setting, methods based on sample consistently exhibit stronger performance than those relying on model-level aggregation. Specifically, approaches that directly optimize a mixture of source and target samples—such as WADN, MADA, OTQMS, and UOWQ—tend to outperform model-based techniques like H-ensemble and MCW, which combine predictions or parameters from multiple pretrained models. A plausible explanation is that sample-based strategies allow the target model to interact directly with source data during training, enabling more flexible adaptation to target-specific distributions, whereas model-based methods are constrained by fixed source representations.
Unified Optimization Outperforms Isolated Optimization. Among sample-based approaches, UOWQ further improves upon existing baselines by jointly optimizing both the transfer quantities and the corresponding weights, while also deriving transfer weights that are explicitly dependent on the transfer quantities. In contrast, OTQMS focuses solely on optimizing transfer quantities, whereas WADN primarily adjusts weighting coefficients.
Gains over Strong Baselines in Challenging Domains. Beyond average accuracy improvements, UOWQ consistently outperforms strong baselines in the most challenging domains. For example, in the DomainNet experiments, UOWQ improves over the previous best results in the Quickdraw target domain by 3.0% absolute accuracy (36.8% vs. 33.8%) and in the Infograph target domain by 1.4% (35.2% vs. 33.8%). These gains in visually divergent domains demonstrate that UOWQ’s sample-weighting strategy is particularly effective in mitigating domain shift. These results highlight the robustness of UOWQ under severe domain shifts, where knowledge transfer is particularly challenging.
VI-D Multi-Task Learning Performance
The results for the multi-task learning setting are summarized in Table III. We make the following observations.
Extension of Weighting Theory to Multi-Task Learning. Although the proposed weighting theory is derived under a multi-source transfer learning formulation, our experiments demonstrate that it generalizes effectively to the multi-task learning setting. This observation validates the effectiveness and generality of the weights introduced by UOWQ. Specifically, UOWQ iteratively optimizes all model parameters and adaptively regulates cross-task influence through weights derived from our theoretical analysis and explicitly related to the transfer quantities.
VI-E Robustness under Different Shot Settings
We further investigate the robustness of UOWQ with respect to the availability of target samples by gradually increasing the number of shots per class from 5 to 100. Fig. 4 reports the results on the Office-Home dataset across all four target domains (Ar, Cl, Pr, and Rw), in comparison with Target-Only and AllSources $\cup$ Target. Across all domains, UOWQ maintains consistently competitive accuracy throughout the entire range of shot settings, exhibiting neither performance collapse in low-shot regimes nor degradation as more target samples are introduced. In contrast to methods that rely heavily on either source aggregation or target-only supervision, UOWQ exhibits stable and monotonic performance trends as the number of shots increases. These observations indicate that the proposed approach adapts smoothly to different data regimes and remains robust under varying levels of target supervision.
(a) DomainNet: $\alpha$ weights
(b) DomainNet: $w$ weights
(c) Office-Home: $\alpha$ weights
(d) Office-Home: $w$ weights
(a) DomainNet: $\alpha$ weights
(b) DomainNet: $w$ weights
(c) Office-Home: $\alpha$ weights
(d) Office-Home: $w$ weights
Figure 6: Domain preference visualization on DomainNet and Office-Home trained with half of the source dataset.
VI-F Weight Visualization
To analyze the domain preference learned by UOWQ, we visualize the optimal coefficients $\alpha^{*}$ (as defined in (23)) together with the base weights $w^{*}$(as defined in  (25)) in Fig. 5, where each row corresponds to a target domain and each column corresponds to a source domain. On DomainNet, when Clipart and Painting are treated as target domains, the Real domain receives the largest $\alpha^{*}$ and $w^{*}$ coefficients. In addition, the Quickdraw domain receives relatively small weights for any target domain. These observed domain correlations are consistent with previous findings reported in [39]. In Office-Home, the Real-World domain obtains the highest $\alpha^{*}$ for Art, Clipart, and Product targets, which is consistent with previous empirical observations in [26]. Overall, these results suggest that the quadratic optimization naturally assigns larger transfer coefficients to empirically more transferable source domains. ă
Moreover, Fig. 6 visualizes the learned domain preferences when only half of the source samples are used. Compared with Fig. 5, the relative patterns of the normalized coefficients $\alpha$ remain largely unchanged, indicating that the inferred domain similarity is stable. In contrast, the absolute source weights $w$ become noticeably larger across most source–target pairs. This behavior is consistent with our theory: when the source sample size $N_{i}$ is reduced, the source effect weakens, and the optimal solution compensates by assigning larger source weights to maintain a favorable bias–variance trade-off. From the (25), since $w^{*}_{i}=\frac{s^{*}{\alpha}^{*}_{i}}{N_{i}}$ explicitly depends on $N_{i}$, decreasing $N_{i}$ naturally leads to increased $w_{i}$. Overall, the comparison between Fig. 5 and Fig. 6 shows that UOWQ adaptively adjusts source weights according to transfer quantities rather than relying on fixed weighting schemes.
VI-G Computational Efficiency.
We further compare the computational cost of different methods on the DomainNet benchmark, and report the wall-clock training time in Table IV. Compared with MADA, which requires substantially longer training time across all target domains, both AllSources $\cup$ Target and UOWQ exhibit significantly improved efficiency. Notably, UOWQ achieves comparable training time to AllSources $\cup$ Target, with only a marginal overhead introduced by the additional optimization procedure. Across all domains, the average training time of UOWQ (10:00:32) remains close to that of AllSources $\cup$ Target (09:50:10), and is substantially lower than MADA (17:07:55). These results indicate that the proposed method introduces minimal computational overhead while providing consistent performance gains, demonstrating a favorable trade-off between effectiveness and efficiency.
VI-H Static vs. Dynamic Weighting.
VI-I Compatibility to LoRA-Based Training
We further examine the compatibility of UOWQ with LoRA. Specifically, we apply UOWQ to multi-source transfer on the Office-Home dataset using the ViT-B backbone. Each domain is treated as a distinct source task, and LoRA modules with rank 8 are inserted into the projection layers of transformer blocks. As summarized in Table VI, UOWQ consistently improves accuracy over baselines, confirming that the proposed theoretical principles remain effective under LoRA-based adaptation.
VII Conclusion
@@
In this work, we propose UOWQ, a theoretical framework for the unified optimization of source weights and transfer quantities in multi-source transfer learning. By formulating the training process as a parameter estimation problem and conducting an asymptotic analysis of a K-L generalization error measure, we derive a principled method for computing the optimal source weights and transfer quantities. A notable theoretical insight is that, when the source weights are jointly optimized, the optimal strategy is to use all available source samples. This is because enlarging the transfer quantities reduces estimation variance, while adaptive weighting suppresses the bias caused by heterogeneous domains. Building on these theoretical results, we further develop practical algorithms that assign source weights and update them dynamically during training for both multi-source transfer learning and multi-task learning scenarios. Extensive experiments on real-world datasets demonstrate that UOWQ consistently improves performance over state-of-the-art baselines, particularly in few-shot and cross-domain settings. These results validate both the theoretical foundations and the practical advantages of our approach.
Appendix A Justification of the KL-Based Measure
$-\sum_{z\in\mathcal{Z},y\in\mathcal{Y}}P(z,y)\log{\hat{P}(z)},$ (26)
where the second term of (A), i.e.,
is the standard definition of generalization error in a classification task. Furthermore, the first term of (A) depends solely on the true data distribution and is constant. Though the third term involves $\hat{P}(z)$, $\hat{P}(z)$ is typically treated as a fixed marginal distribution decided by the training data rather than being learned in the model training. As a result, the third term is usually not parameterized by model parameter $\hat{\theta}$, and thus does not contribute to the optimization objective. In brief, the first term and the third term will not affect our optimization. Therefore, the K-L divergence formulation remains consistent with the standard definition of generalization error.
Appendix B Proof of Theorem 2
@@
Lemma 6.
[…]
$o\left(\mathbb{E}\left[\left(\hat{\theta}-\theta_{0}\right)^{2}\right]\right)+o(\frac{1}{N_{0}}).$ (27)
Proof.
$D\left(P_{X;\theta_{0}}\middle\|P_{X;\hat{\theta}(X^{N_{0}})}\right)$
We denote $\delta$ as a small constant, and we can rewrite (1) as
$\left(\frac{1}{2}\sum_{x\in X}\frac{\left(P_{X;\hat{\theta}(X^{N_{0}})}(x)-P_{X;\theta_{0}}(x)\right)^{2}}{P_{X;\theta_{0}}(x)}+o(|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2})\right)$
$+\sum_{\left\{X^{N_{0}}:|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2}\geq\delta\right\}}P_{X^{n};\theta_{0}}(X^{N_{0}})\cdot D\left(P_{X;\theta_{0}}\middle\|P_{X;\hat{\theta}(X^{N_{0}})}\right)$
$+\sum_{\left\{X^{N_{0}}:|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2}\geq\delta\right\}}P_{X^{n};\theta_{0}}(X^{N_{0}})\cdot D\left(P_{X;\theta_{0}}\middle\|P_{X;\hat{\theta}(X^{N_{0}})}\right)$
Definition 7.
[…]
We denote $\hat{P}_{X^{N_{0}}}$ as the empirical distribution of $X^{N_{0}}$. Applying Sanov’s Theorem to (B), we can know that
Then, we aim to establish a connection between (31) and $|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2}$. From (B), we can know that the $D\left(\hat{P}_{X^{N_{0}}}\middle\|P_{X;\theta_{0}}\right)$ in (31) can be transformed to
$D\left(\hat{P}_{X^{N_{0}}}\middle\|P_{X;\theta_{0}}\right)$
From the characteristics of MLE, we can know that
$=0$
$+O(|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2}),$ (33)
which can be transformed to
Using the Cauchy-Schwarz inequality, we can obtain
where
Combining with (B), (B), and (B), the inequality (B) can be transformed to
$D\left(\hat{P}_{X^{N_{0}}}\middle\|P_{X;\theta_{0}}\right)$
$+o(|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2})$
Combining (31) and (B), we can know that
For the first term in (B), the negligible term $o(|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2})$ is of lower order than the main term and can therefore be ignored. As for the second term in (B), its probability is bounded by $O(e^{\frac{-N_{0}J(\theta_{0})|\hat{\theta}(X^{N_{0}})-\theta_{0}|^{2}}{2}})$ according to (38), which decays exponentially with $N_{0}$. Consequently, this probability is $o(1/N_{0})$, since exponential decay in $N_{0}$ dominates any polynomial rate. Moreover, under the compactness assumption on the parameter space stated in Section III-C, the K-L divergence is uniformly bounded, implying that the entire second term is $o(1/N_{0})$. By transferring (B), we can get
$=\frac{1}{2}\mathbb{E}\left[\sum_{x\in X}\frac{\left(\frac{\partial P_{X;\theta_{0}}(x)}{\partial\theta_{0}}(\hat{\theta}(X^{N_{0}})-\theta_{0})\right)^{2}}{P_{X;\theta_{0}}(x)}\right]$
$+o\left(\mathbb{E}\left[\left(\hat{\theta}-\theta_{0}\right)^{2}\right]\right)+o{(\frac{1}{N_{0}})}.$ (39)
We then transform (B) with (B)
Combining (B), (B), we can get
[…]
$o\left(\mathbb{E}\left[\left(\hat{\theta}-\theta_{0}\right)^{2}\right]\right)+o{(\frac{1}{N_{0}})}.$ (41)
We denote $E_{\hat{\theta}}$ as the expectation of $\hat{\theta}$. Then, we can make a decomposition of the K-L measure, i.e.,
[…]
$=\frac{1}{2}{J}(\theta_{0})\left(\mathbb{E}\left[\left(\hat{\theta}-E_{\hat{\theta}}\right)^{2}\right]+\mathbb{E}\left[\left(E_{\hat{\theta}}-\theta_{0}\right)^{2}\right]\right)$
$+o\left(\mathbb{E}\left[\left(\hat{\theta}-E_{\hat{\theta}}\right)^{2}\right]+\mathbb{E}\left[\left(E_{\hat{\theta}}-\theta_{0}\right)^{2}\right]\right)+o(\frac{1}{N_{0}}).$ (42)
[…]
@@
Combining with (58), we can transform the numerator of the right-hand side of (59) as follows:
@@
Applying the Central Limit Theorem to (B), we can get
$\sqrt{N_{0}+w_{1}n_{1}}L_{n}^{{}^{\prime}}(E_{\hat{\theta}})\xrightarrow{a.s.}$ (61)
@@
Similarly to (B) and (B), we can prove
$\mathbb{E}_{\theta_{1}}\left[\left(\frac{\partial\log{P_{X;E_{\hat{\theta}}}(x)}}{\partial\theta}\right)^{2}\right]=J(\theta_{1})+O(\frac{1}{\sqrt{N_{0}}}),$ (65)
and
$\mathbb{E}_{\theta_{1}}\left[\left(\frac{\partial\log{P_{X;E_{\hat{\theta}}}(x)}}{\partial\theta}\right)\right]^{2}=O(\frac{1}{N_{0}}).$ (66)
By substituting (B), (B), (65), and (66) into (61), we obtain
@@
Substituting (47) and (71) into (B), we have
@@
Similarly, by taking the derivative of (B) with respect to $n_{1}$, we observe that the derivative is
@@
Appendix C Proof of Proposition 3